Please show work I am not getting the correct answers

1. Hydrogen peroxide, H2O2, decomposes spontaneously through the following reaction.

2H2O2(aq) 2H2O(g) + O2(g)

The kinetics are second order, and the half‐life is 17.3 min when a 1.75 M solution of H2O2 is made.

What is the concentration of the solution after 24.8 min?

2. The decomposition of N2O5 occurs by the following reaction:

2N2O5(g) 4NO2(g) + O2(g)

The rate constant is experimentally determined to be 3.4 x 10‐4 s‐1. A sample is sealed in a flask and

after 6 min the concentration of N2O5 is 2.36 M. Calculate the initial concentration of the N2O5.
Answer #1

1)
Given:
Half life = 17.3 min
use relation between rate constant and half life of 2nd order reaction

k = 1/([H2O2]o*half life)
= 1/(1.75*17.3)
= 3.303*10^-2 M-1.min-1
we have:
[H2O2]o = 1.75 M
t = 24.8 min
k = 3.303*10^-2 M-1.min-1


use integrated rate law for 2nd order reaction
1/[H2O2] = 1/[H2O2]o + k*t
1/[H2O2] = 1/(1.75) + 3.303*10^-2*24.8
1/[H2O2] = 0.5714 + 3.303*10^-2*24.8
1/[H2O2] = 1.3906
[H2O2] = 0.719 M
Answer: 0.719 M

Only 1 question at a time please